# Instruction Tuning for Large Language Models: A Survey

### v7

@@ 8.5 Smaller Instruction Dataset
IT requires a substantial amount of specialized instruction data for training. Zhou et al. (2023a) hypothesized that the pre-trained LLM only has to learn the style or format to interact with users and proposed LIMA that achieves strong performance by fine-tuning an LLM on only 1,000 carefully selected training examples.
@@ 8.6 Evaluating Instruction Tuning Datasets
The performance of IT model highly depends on the IT datasets. However, there lacks of evaluations for these IT datasets from open-ended and subjective aspects.
To address this issue, Wang et al. (2023d) performs dataset evaluation by fine-tuning the LLaMa model Touvron et al. (2023b) on various of open IT datasets and measure different fine-tuned models through both automatic and human evaluations. An additional model is trained on the combination of IT datasets. For the results, Wang et al. (2023d) showed that there is not a single best IT dataset across all tasks, while by manually combining datasets it can achieve the best overall performance. Besides, Wang et al. (2023d) pointed out that though IT can bring large benefits on LLMs at all sizes, smaller models and models with a high base quality benefit most from IT. For human evaluations, Wang et al. (2023d) a larger model is more likely to gain a higher acceptability score.
8.7 Superficial Alignment
To answer these questions, Kung and Peng (2023) delves into the analysis of how models make use of instructions during IT by comparing the tuning when provided with altered instructions versus the original instructions.
Specifically, Kung and Peng (2023) creates simplified task definitions that remove all semantic components, leaving only the output information. In addition, Kung and Peng (2023) also incorporates delusive examples that contain incorrect input-output mapping. Surprisingly, the experiments show that models trained on these simplified task definitions or delusive examples can achieve comparable performance to the ones trained on the original instructions and examples. Moreover, the paper also introduces a baseline for the classification task with zero-shot, which achieves similar performance to IT in low-resource settings.
Similar to the findings of Kung and Peng (2023), several subsequent studies Zhou et al. (2023a); Lin et al. (2023a) reached the same conclusion: the observed performance improvements in current IT models are often due to superficial alignment. This means the models excel at recognizing superficial alignment, such as mastering output formats and making educated guesses, rather than truly understanding and learning the underlying tasks.
@@ 9 Conclusion
This work surveys recent advances in the fast growing field of instruction tuning, which can also be referred to as supervised fine-tuning (SFT). We make a systematic review of the literature, including the general methodology of IT, the construction of IT datasets, the training of IT models, IT’s applications to different modalities, domains and application. We also review analysis on IT models to discover both their advantages and potential pitfalls. We hope this work will act as a stimulus to motivate further endeavors to address the deficiencies of current IT models.
@@ Appendix A Datasets
    17 https://github.com/Instruction-Tuning-with-GPT-4/GPT-4-LLM
@@
    20 https://github.com/LAION-AI/Open-Assistant

### v9

@@ 8.5 Smaller Instruction Dataset
SFT requires a substantial amount of specialized instruction data for training. Zhou et al. (2023a) hypothesized that the pre-trained LLM only has to learn the style or format to interact with users and proposed LIMA that achieves strong performance by fine-tuning an LLM on only 1,000 carefully selected training examples.
@@ 8.6 Evaluating Instruction Tuning Datasets
The performance of SFT model highly depends on the SFT datasets. However, there lacks of evaluations for these SFT datasets from open-ended and subjective aspects.
To address this issue, Wang et al. (2023e) performs dataset evaluation by fine-tuning the LLaMa model Touvron et al. (2023b) on various of open SFT datasets and measure different fine-tuned models through both automatic and human evaluations. An additional model is trained on the combination of SFT datasets. For the results, Wang et al. (2023e) showed that there is not a single best SFT dataset across all tasks, while by manually combining datasets it can achieve the best overall performance. Besides, Wang et al. (2023e) pointed out that though SFT can bring large benefits on LLMs at all sizes, smaller models and models with a high base quality benefit most from SFT. For human evaluations, Wang et al. (2023e) a larger model is more likely to gain a higher acceptability score.
@@ 8.7 Proprietary LLMs Imitation
9 The Role of Instruction Fine-tuning
Instruction fine-tuning (IF), also known as supervised fine-tuning (SFT), is a conventional alignment approach that trains models on example prompts paired with corresponding responses to ensure the model’s output aligns with user instructions and intended goals. More recently, some reinforcement learning (RL) based methods Wang et al. (2024), e.g., reinforcement learning from human feedback (RLHF) Ouyang et al. (2022), direct preference optimization (DPO) Rafailov et al. (2023), and group relative policy optimization (GRPO) Shao et al. (2024), and various prompt engineering strategies have emerged as alternatives or complements. Thus, in this section, we will review each method’s role in aligning LLMs, and examine whether SFT remains necessary. Further more, we also consider the risk of superficial alignment, i.e. alignment that changes only the model’s surface behavior (tone, style) without imparting deeper understanding.
9.1 SFT Compared with Other Alignment Methods
Below, we begin by outlining three widely used alignment approaches, RLHF, DPO, and prompt engineering, highlighting their strengths and weaknesses. Then, we explain why SFT remains an essential component of contemporary alignment pipelines.
9.1.1 Reinforcement Learning from Human Feedback (RLHF)
RLHF is the dominant alignment paradigm, and typically proceeds in three phases: (1) supervised fine-tuning (SFT) on human large amounts of instruction-answer pairs, (2) training a reward model on human-ranked responses, and (3) using policy optimization (e.g. PPO schulman2017proximal) to maximize the reward model’s feedback Chen et al. (2025b). This pipeline can deeply adjust model behavior to complex user preferences. RLHF has enabled remarkable capabilities (e.g. nuanced help, factuality), but at high cost and complexity. It requires extensive human data, careful RL tuning, and often suffers stability issues and “reward hacking” (the model finds loopholes in the reward model) Xiao et al. (2024); Wang et al. (2024). Because RLHF optimization is resource-intensive and sensitive to hyper parameters, simpler alternatives have been sought.
Advantages.
The key strength of RLHF lies in its ability to guide models toward high-level objectives, such as helpfulness and safety, that are not explicitly encoded in the training data, demonstrating strong empirical performance in aligning models with user intent when well-tuned Wang et al. (2024); Chen et al. (2025b).
Limitations.
RLHF’s complexity is a downside. It typically requires starting from an SFT-trained model, i.e., a model that already follows instructions to some degree, because training RL from a raw base model is difficult Trivedi et al. (2025). The multi-stage pipeline (SFT, reward model, and PPO) is time-consuming and brittle Chen et al. (2025b). In practice, researchers and practitioners often still perform an initial instruction fine-tuning, even when using RLHF, to establish a reasonable base policy. Moreover, RLHF can introduce “alignment tax” (performance drop on some tasks) and can fail to generalize if the reward model is mis-specified Xiao et al. (2024).
9.1.2 Direct Preference Optimization (DPO)
Direct Preference Optimization (DPO) Rafailov et al. (2023) is a recently proposed RL-free alignment method that directly fine-tunes on preference pairs. Instead of learning a separate reward model and running RL, DPO casts alignment as a supervised objective: for each prompt and pair of outputs (preferred vs. dispreferred), it adjusts the model’s logits to increase the probability of the preferred output. DPO’s loss is equivalent to a Bradley–Terry pairwise classification (a logit-ratio objective) to bypass policy-gradient RL entirely.
Advantages.
Because DPO fits into a standard maximum-likelihood fine-tuning framework, it is far simpler and more stable than PPO-based RLHF Xu et al. (2024a); Xiao et al. (2024); Wang et al. (2024) . Studies report that DPO matches or exceeds RLHF performance on tasks like summarization or helpfulness with fewer preference examples. Compared to RLHF, DPO has been shown to be stable, performant, and computationally lightweight in various applications. It does not require expensive RL infrastructure or tuning of PPO hyper-parameters, making it reproducible and easier to deploy. Practitioners, e.g. OpenAI  [^20] https://cookbook.openai.com/examples/fine_tuning_direct_preference_optimization_guide and DeepSeek Shao et al. (2024) note that starting DPO from an already fine-tuned model improves results, using SFT to establish a robust initial policy.
Limitations.
DPO still depends on quality preference data, and like any offline method it can suffer if the data distribution shifts. Recent analyses have identified issues with DPO: because it employs an implicit reward tied to the policy, it can bias the model toward out-of-distribution outputs and even degrade generalization Xiao et al. (2024); Wang et al. (2024). Variants, e.g. KL-constrained or semantics-aware DPO, are being developed to mitigate these problems, but it remains true that DPO typically benefits from starting with a good initial model. In practice, most implementations still use an SFT-tuned model before running DPO, echoing the RLHF pipeline. Thus, while DPO simplifies alignment, it has not eliminated the need for supervised tuning in many cases.
9.1.3 Prompt Engineering (In-Context Learning)
Prompt engineering aligns model behavior without any fine-tuning. Instead, it leverages the model’s existing capabilities by crafting prompts, including instructions, few-shot examples, or chain-of-thought cues, to elicit desired outputs. Recent work treats prompt design itself as an optimization problem: one can optimize a prompt string or learn soft prompts to maximize human-aligned metrics Trivedi et al. (2025). Importantly, prompt-based methods assume no weight updates, which means that they work without any post training.
Advantages.
The biggest benefit is that no retraining is required. Prompt optimization can effectively align LLMs even when parameter fine-tuning is not feasible Trivedi et al. (2025). This is appealing for large models with fixed parameters (APIs or frozen on-device models). Some experiments have shown that with well-designed prompts, base LLMs can achieve high performance on instruction-following tasks. For example, Lin et al. (2023b) introduced URIAL, a method that uses only a few stylistic examples in-context (plus a system prompt) to steer the model, and found it matched or even surpassed fully-tuned models in many benchmarks. This suggests that clever prompt engineering alone can yield strong alignment in some cases Wang et al. (2023c); Sun et al. (2023d, b); Wang et al. (2023b); Sun et al. (2023c). Prompting has obvious speed and convenience advantages: it requires no training data or compute, and can be iterated quickly by users.
Limitations.
Prompt-based alignment has inherent limits. The model’s context window bounds how much instruction or example content can be provided so that very complex tasks may simply not fit. More importantly, prompt methods generally induce superficial compliance rather than truly altering the model’s knowledge. They leverage already-encoded patterns in the model, but cannot add new capabilities or correct deep misunderstandings. In practice, prompt engineering often produces brittle behaviors: slight rephrasing can break performance, and malicious users can “jailbreak” around prompts to elicit bad outputs. For instance, Chen et al. (2025a) note that although prompt-based ICL can align a model to some extent, it does so mainly by inserting stylistic cues and does not fundamentally change the model’s reasoning process. In short, prompt engineering can quickly achieve surface-level alignment (tone, disclaimers, formatting), but cannot replace weight tuning for deep or novel tasks, and for complex reasoning, mathematics, or new knowledge integration typically require additional fine-tuning.
9.1.4 The Continued Necessity of SFT
Given these techniques, a key question is whether instruction fine-tuning (IF) or supervised fine-tuning (SFT) remains necessary in modern pipelines. Empirical evidence suggests it does. Both RLHF and DPO pipelines almost universally incorporate an initial SFT stage. In RLHF this is explicit, which SFT usually serves as the first phase. For DPO, while the final optimization is simpler, practitioners generally first fine-tune on good example responses to establish a robust initial policy, which stabilizes subsequent DPO refinement. The OpenAI alignment guide explicitly recommends performing supervised fine-tuning on a subset of preferred responses before DPO to improve alignment and convergence [^21] https://cookbook.openai.com/examples/fine_tuning_direct_preference_optimization_guide. In other words, even with DPO, a round of instruction tuning yields better outcomes. On the other hand, prompt-based methods show that in principle one can align some models without any fine-tuning Wang et al. (2023c); Sun et al. (2023d, b); Wang et al. (2023b); Sun et al. (2023c). Other work Lin et al. (2023a) demonstrates that you can in effect “unlock” base LLMs by providing few-shot prompts, achieving performance close to tuned models without SFT. However, these tuning-free methods tend to rely on pre-existing capabilities. If a base model genuinely lacks a skill, such as solving a type of math problem it was never pre-trained on, no prompt will fix it, and only updating weights can. Moreover, some recent studies Parthasarathy et al. (2024) find that for reasoning and knowledge tasks, performance continues to scale with more fine-tuning data, suggesting base models do improve under instruction tuning. In summary, prompt methods can sometimes obviate SFT for shallow alignment, but for robust alignment pipelines or domain specific alignment (e.g., Medicinal Chemistry), supervised fine-tuning is still regarded as essential groundwork. New research even explores hybrid tricks, such as “instruction residuals” from an older model added to a new base, to avoid re-training, but these rely on existing tuned models as sources. The prevailing practice remains: use SFT to teach the model the format and style of responses, then refine preferences via RLHF or DPO.
9.2 Superficial Alignment
To answer these questions, Kung and Peng (2023) delves into the analysis of how models make use of instructions during SFT by comparing the tuning when provided with altered instructions versus the original instructions.
Specifically, Kung and Peng (2023) creates simplified task definitions that remove all semantic components, leaving only the output information. In addition, Kung and Peng (2023) also incorporates delusive examples that contain incorrect input-output mapping. Surprisingly, the experiments show that models trained on these simplified task definitions or delusive examples can achieve comparable performance to the ones trained on the original instructions and examples. Moreover, the paper also introduces a baseline for the classification task with zero-shot, which achieves similar performance to SFT in low-resource settings.
Similar to the findings of Kung and Peng (2023), several subsequent studies Zhou et al. (2023a); Lin et al. (2023a) reached the same conclusion: the observed performance improvements in current SFT models are often due to superficial alignment. This means the models excel at recognizing superficial alignment, such as mastering output formats and making educated guesses, rather than truly understanding and learning the underlying tasks.
@@ 10 Conclusion
This work surveys recent advances in the fast growing field of instruction tuning, which can also be referred to as supervised fine-tuning (SFT). We make a systematic review of the literature, including the general methodology of SFT, the construction of SFT datasets, the training of SFT models, SFT’s applications to different modalities, domains and application. We also review analysis on SFT models to discover both their advantages and potential pitfalls. We hope this work will act as a stimulus to motivate further endeavors to address the deficiencies of current SFT models.
@@ Appendix A Datasets
    17 https://github.com/LAION-AI/Open-Assistant
@@
    20 https://github.com/Instruction-Tuning-with-GPT-4/GPT-4-LLM
@@
  • 30
    30 https://github.com/openai/prm800k
  • 31
    31 https://github.com/GAIR-NLP/O1-Journey
  • 32
    32 https://github.com/MARIO-Math-Reasoning/MARIO
  • 33
    33 https://github.com/deepseek-ai/DeepSeek-Math
  • 34
    34 https://github.com/XylonFu/MathScale
  • 35
    35 https://github.com/pipilurj/G-LLaVA